OpenStudy (anonymous):

An electron in an excited state of a hydrogen atom emits two photons in succession, the first at 2165 nm and the second at 97.20 nm, to return to the ground state (n=1). For a given transition, the wavelength of the emitted photon corresponds to the difference in energy between the two energy levels.What were the principal quantum numbers of the initial and intermediate excited states involved?

OpenStudy (aaronq):

Use the rydberg formula: \(\dfrac{1}{\lambda}=E=R(\dfrac{1}{n^2_f}-\dfrac{1}{n^2_i})\) R=\(1.097*10^7 \;m^{-1}\)